A company needs an increase in working capital of $50,000 in a project that will last 4 years. The companys tax rate is 30% 4 and its after-tax discount rate is 8%. Click here to view Exhibit 13B-1 to determine the appropriate discount factor(s) using table. The present value of the release of the working capital at the end of the project is closest to: 02:-37.10 Multiple Choice $36,750 $15,000 $25,726 < Prev 40.10 Next>
5: Rhoads Corporation is considering a capital budgeting project that would require an investment of $160.000 in equipment with a 4-year expected life and zero salvage value. Annua, incremental sales will be $460,000 and annual İncremental cash operating expenses vil be $330,000 The companys income tax rate is 30% and the after-tax discount rate is 15, The company uses straight ine depreciation on all equipment; the annual depreciation expense will be $40.000. Assume cash fows occur at the end of the year except for the initial investments. The company takes income taxes into account in its )Capital buclgeting. 08-0 48 Click here to view Exhibit 138-1 to determine the appropriate discount factors) using table. The net present value of the project is closest to:
5 Multiple Choice 02:36:24 O $178.252 $252,000 $97,040 $134,168
6: Fontana Corporation is considering a capital budgeting project thot would require investing $240,000 in equipment with an expected life of 4 years and zero salvage value. The annual incremental sales would be $640,000 and the annual incremental cash operating expenses would be $440,000. The companys income tax rte is 30%, The company uses straight line depreciation on all equipment 02:36 03 The total cash flow net of income taxes in year 2 is: Multiple Choice $158,000 $200,00o
Multiple Choice 6 $158,000 ( 02:35:49) $200,000 $88,000 $140,000
Answer #1
Ans (4) Present value of release of working capital at the end of project is closest to $36750
Working Capital release in year 4 $           50,000.00
P.V @ 8% for 4 years 0.735
P.V $           36,750.00
Rhoads Corporation
Year 0 1 2 3 4
Investment $     (1,60,000.00)
Revenue=(A) $        4,60,000.00 $     4,60,000.00 $        4,60,000.00 $      4,60,000.00
Expenses=(B) $        3,30,000.00 $     3,30,000.00 $        3,30,000.00 $      3,30,000.00
Depreciation=($160000/4)=(C) $           40,000.00 $         40,000.00 $            40,000.00 $          40,000.00
Income before tax=(D)=(A)-(B)-(C ) $           90,000.00 $         90,000.00 $            90,000.00 $          90,000.00
Taxes=(E )=(D)*30% $           27,000.00 $         27,000.00 $            27,000.00 $          27,000.00
Profit after tax=(F )=(D )-(E ) $           63,000.00 $         63,000.00 $            63,000.00 $          63,000.00
Add: Depreciation=(G) $           40,000.00 $         40,000.00 $            40,000.00 $          40,000.00
Cash Inflow=(H)=(F)+(G) $        1,03,000.00 $     1,03,000.00 $        1,03,000.00 $      1,03,000.00
Cash Inflow=(I) $     (1,60,000.00) $        1,03,000.00 $     1,03,000.00 $        1,03,000.00 $      1,03,000.00
P.V Factor @15%=(J) 1 0.87 0.756 0.658 0.572
Present Value=(I)*(J) $     (1,60,000.00) $           89,610.00 $         77,868.00 $            67,774.00 $          58,916.00
Ans 5 NPV $       1,34,168.00
Ans 6) $158000
Fontana Corporation
Year 0 1 2 3 4
Investment $     (2,40,000.00)
Revenue=(A) $        6,40,000.00 $     6,40,000.00 $        6,40,000.00 $      6,40,000.00
Expenses=(B) $        4,40,000.00 $     4,40,000.00 $        4,40,000.00 $      4,40,000.00
Depreciation=($240000/4)=(C) $           60,000.00 $         60,000.00 $            60,000.00 $          60,000.00
Income before tax=(D)=(A)-(B)-(C ) $        1,40,000.00 $     1,40,000.00 $        1,40,000.00 $      1,40,000.00
Taxes=(E )=(D)*30% $           42,000.00 $         42,000.00 $            42,000.00 $          42,000.00
Profit after tax=(F )=(D )-(E ) $           98,000.00 $         98,000.00 $            98,000.00 $          98,000.00
Add: Depreciation=(G) $           60,000.00 $         60,000.00 $            60,000.00 $          40,000.00
Cash Inflow=(H)=(F)+(G) $        1,58,000.00 $ 1,58,000.00 $        1,58,000.00 $      1,38,000.00